# Emergence of two inertial sub-ranges in solar wind turbulence: dependence on heliospheric distance and solar activity

### v1

@@ 4.1 Observations during Solar Minimum
Figure 3: Magnetic power spectral trace of FSW intervals plotted for nearly equidistant frequencies at varying heliospheric distances $R$ during solar minima. Top row: Helios data, bottom row: Ulysses data. In all panels, the distance of the interval from the sun is indicated. Vertical lines indicate the $f^{-1}$ break (dashed), the newly observed break $f_{b}$ (dot-dashed, separating the light and deep blue areas jointly forming the traditional inertial range), and the ion-scale break (Ulysses only). In each range, a power-law fit is shown (colored lines) along with the corresponding spectral exponent.
Figure 4: Kurtosis $K(\tau)$ of magnetic field fluctuations for several intervals of FSW during periods of solar minima. Top panels: Helios data (year 1976) in the inner heliosphere from a sustained coronal hole. Bottom panels: Ulysses data (years 1995-1996) in the outer heliosphere at varying distances and latitudes (the distance of each interval is indicated and associated to a given color). Power-law fits and the corresponding scaling exponents are indicated. Vertical lines indicate the observed break, $\tau_{K}$ (solid lines), and the timescale corresponding to the spectral break, $1/f_{b}$ (dashed).
@@
During a period of solar minimum in 1976, using data from Helios spacecraft, we study FSW streams in the inner heliosphere (at $0.3,0.41,0.65,$ and $0.98$ au) from a sustained coronal hole near the ecliptic plane. Beyond $1$ au, FSW streams are studied using Ulysses data collected during the 1995-1996 solar minimum, at varying heliospheric distances (at $1.44,2.1,2.75,$ and $3.16$ au), which were also measured at different latitudes. In Fig. 3, we have drawn the magnetic power spectral traces, estimated using nearly equispaced frequencies in logarithmic scales. Top panels refer to Helios intervals, while bottom panels to Ulysses. As typically observed in the Alfvénic solar wind, at low frequencies we can identify a large-scale, energy-containing range (white background in the figure), where the power decays as $\sim f^{-1}$. Fitted power laws and the corresponding scaling exponents are shown as green lines. A break identifies a clear change in the power-law scaling exponent, as indicated by vertical dashed lines. Such break can be associated with the correlation scale of the turbulence. The low-frequency range is clearly visible in Helios data, while it is only indicatively present in the Ulysses intervals. This is consistent with the well-known shift of the correlation scale towards lower frequency with increasing $R$ in the solar wind (Davis et al., 2023). The $f^{-1}$ range is followed by the usual inertial range of turbulence, where the spectrum roughly follows an $f^{-5/3}$ power law dependence (Bruno & Carbone, 2005, 2013). However, a more accurate inspection shows that a further break emerges within such range, indicated by the vertical dot-dashed lines separating the light and deeper blue shaded areas in Fig. 3. Although the dynamical range of frequencies is relatively short, for intervals other than that at $3.16$ au, it is possible to identify two different sub-ranges with different power laws as demonstrated by the red and blue lines, with the associated scaling exponents indicated nearby. In the lower-frequency range (light blue background), the spectral index approaches -3/2, whereas at larger frequencies (deep blue background), the spectra show a transition to a -5/3 spectral index usually observed in non-Alfvénic solar wind (Bruno & Carbone, 2005, 2013; Alexandrova et al., 2009; D’Amicis et al., 2018). In isotropic turbulence, whereas an $f^{-5/3}$ scaling often represents an energy cascade by eddy fragmentation in strong turbulence, $f^{-3/2}$ scaling can possibly be explained by an energy cascade through the sporadic interaction of Alfvénic wave packets in MHD turbulence (Kolmogorov, 1941; Iroshnikov, 1963; Kraichnan, 1965). However, $-5/3$ and $-3/2$ power laws can also be obtained under various circumstances if anisotropy is taken into account (Goldreich & Sridhar, 1995, 1997; Boldyrev, 2006; Chandran et al., 2015). Irrespective of the true nature of energy cascade, a single power law is often assumed for the magnetic power spectra in the frequency range $10^{-4}-10^{-1}$ Hz (Bruno & Carbone, 2005, 2013), although a few studies have found variation in the power law exponents in the inertial range of magnetic power spectra (Wicks et al., 2011; Sioulas et al., 2023b) as well as the scaling of higher order structure functions (Wu et al., 2022; Sorriso-Valvo et al., 2023). In our study, the co-existence of the two sub-regimes (with $-3/2$ and $-5/3$ spectral indices) within the turbulence spectra of FSW has been consistently observed at various heliospheric distances both in the inner as well as the outer heliosphere. The break scale between those two sub-ranges, $f_{b}$, appears to shift towards lower frequencies (approaching the correlation scale) with increasing heliospheric distance. This is consistent with the fact that a ${-3/2}$ scaling has been observed for solar wind close to the sun, whereas a steeper ${-5/3}$ power law is obtained at and beyond $1$ au (Chen et al., 2020; Shi, C. et al., 2021; Sioulas et al., 2023a). Finally, in the Ulysses intervals, the ion-scale breaks are visible, separating the MHD range from the sub-ion range, where Hall effects and other kinetic effects start to affect the cascade (white background) (Banerjee & Galtier, 2016; Halder et al., 2023). Such break is usually observed at frequencies $\sim 10^{-1}$ Hz, which is the upper cut-off for the MHD range. However, similar breaks do not turn up in the Helios intervals, due to the low cadence of the data used here.
@@
To further investigate on the sub-inertial range spectral break, $f_{b}$, we study the kurtosis $K(\tau)$ for all of the eight FSW intervals. The scaling of $K(\tau)$ defined in Section 3 for Helios and Ulysses data are depicted in Fig. 4 top and bottom panels, respectively, for each $R$. To inspect on the general radial trend of intermittency we have drawn a consolidated plot for the Helios and Ulysses intervals (see Fig. 5). From this figure one can conclude that the value of $K$ at all scales increases with increasing $R$, thus implying higher intermittency with increasing heliospheric distance, in agreement with previous studies (Bruno et al., 2003; Sorriso-Valvo et al., 2023; Sioulas et al., 2023a). At each given distance $R$, $K$ is systematically found to decrease as one moves towards the larger scales. This is consistent with the notion that deviation from Gaussian statistics increases at smaller scales (Frisch, 1995; Sorriso-Valvo et al., 1999). Upon reaching the typical correlation scales of the flow ($\tau\simeq 10^{4}$ s), corresponding to the $f^{-1}$ power law in energy spectrum (see Fig. 3), the kurtosis saturates to a constant value $K\simeq 1.67$, representing a quasi-Gaussian distribution (with a non-zero skewness) of the fluctuations of the magnetic field components (see appendix Section A). Within the inertial range, from the nature of $K(\tau)$ in Fig. 4, a clear signature of broken power law is observed. While two breaks are visible for Ulysses data (with $1$ s resolution), the small-scale break at around $\tau\sim 10$ s is missing for the intervals using Helios magnetic field data with $6$ s resolution. This break, corresponding to a frequency of $\sim 10^{-1}$ Hz, is associated with the transition from the ordinary MHD range to the sub-ion kinetic or Hall MHD regime (Banerjee & Galtier, 2016; Halder et al., 2023). The other break which occurs at a larger $\tau$ (solid vertical lines) is clearly visible for both Helios and Ulysses data. In particular this break scale ($\tau_{K}$) shifts towards larger $\tau$ as $R$ increases. Within the distance range of $0.3-2.75$ au, $\tau_{K}$ is found to increase from $\sim 100$ s to $\sim 1500$ s. It is to be emphasized here that except for certain cases, the appearance of the break $\tau_{K}$ is persistent in the component-wise $K$ scaling as well (see Figs. 10 and 11 in appendix Section B). A detailed list of the break scale $\tau_{K}$ as a function of $R$ is given in Table 2. As it is evident from Fig. 4 and 5, $\tau_{K}$ separates the steeper power law ($K\sim\tau^{-\kappa}$ with $\kappa\simeq 0.37$ averaged over the eight intervals) at smaller scales (dashed lines) from the less steeper one ($\kappa\simeq 0.11$ on average) at large scales (dotted lines), but with an exception. Note that for the Ulysses interval at $R=3.16$ au, $K(\tau)$ reaches the Gaussian regime without going through the large $\tau$ break, suggesting that the turbulence has fully developed that transforms the shallower scaling range at large scale into the steeper power law at smaller scales. We will elucidate this point in the following.
@@
As mentioned in the introduction, similar broken power law behaviour for $K(\tau)$ in FSW has already been observed by Sorriso-Valvo et al. (2023). However, those authors suggested that $\tau_{K}$ might correspond to the break between low-frequency $f^{-1}$ regime to Kolmogorov $f^{-5/3}$ regime in the magnetic power spectra. This was inspired by the fact that $f^{-1}$ regime is exclusively found in FSW intervals and the $f^{-1}$ break also shows nearly similar behaviour to $1/\tau_{K}$ as $R$ changes (Davis et al., 2023). Instead, for all the intervals where the break is observed, it is systematically found in our study that $1/\tau_{K}$ occurs at a higher frequency (roughly by a factor $\sim 10$) than the $f^{-1}$ break scale (see Fig. 3). The inverse of $\tau_{K}$ is typically corresponding to $f_{b}$, although with some consistent small discrepancy that could be due to the different frequency response of Fourier transform and scale-dependent increments (see Fig. 4 where both $\tau_{K}$, solid lines, and $1/f_{b}$, dashed lines, are drawn). The two scaling ranges in the kurtosis therefore approximately correspond to the two inertial sub-ranges observed in the spectrum. Since $PSD$ and kurtosis are related quantities, the observation of a double power law in both supports the robustness of the break, and therefore indicates the emergence of a new characteristic scale in the inertial range that marks the transition from $f^{-3/2}$ to $f^{-5/3}$ regime.
@@
Figure 6: Top: Cross-helicity spectrum $\hat{\sigma_{c}}$ of the four Helios FSW intervals in the inner heliosphere. The light and deep blue shaded regions depict the $-3/2$ and $-5/3$ regimes respectively. The spectral break frequency $f_{b}$ and the frequency associated with the kurtosis break, $\tau_{K}^{-1}$, are indicated by dot-dashed and dashed lines respectively. Bottom: Cross-helicity $\sigma_{c}$ of all the FSW intervals as a function of the heliospheric distance $R$.
@@
Figure 7: FSW during solar minima. Top: Kurtosis scaling break $l_{K}$ versus the heliospheric distance $R$. The same break scale $l_{K}$ normalized by the ion-inertial length scale $d_{i}$, $l_{K}/d_{i}$ (middle) and $l_{K}$ normalized by the correlation length $L_{c}$ (bottom) as a function of $R$. The different colors reproduce the colors in Fig. 3, and in the Ulysses intervals the latitude is indicated. Two power laws were identified in the inner and outer heliosphere, respectively. The fitted power laws and the corresponding parameters are indicated.
Summarizing, from the existence of the two turbulent inertial sub-regimes it is clear that as we move from the larger towards the smaller scales the nature of turbulence also varies. This variation becomes more apparent when we examine the cross-helicity spectrum for the FSW intervals within the inner heliosphere (Fig. 6 top). The same could not be computed for the FSW beyond $1$ au due to the limitation in terms of low plasma data resolution, as mentioned in Section 2. Nevertheless, for all the FSW intervals in the inner heliosphere we see that the $\hat{\sigma_{c}}$ power decreases as we move from larger to smaller scales (see Fig. 6). Thus, with the forward progression of the turbulent cascade, the imbalance between the inward and outward Alfvén modes propagating along the mean magnetic field decreases to a more balanced state. While recent studies have shown the transition from a weak to a strong turbulence regime on moving towards smaller scales (Zhao et al., 2024), a transition from imbalanced ($|z^{+2}|\gg|z^{-2}|$, or vice-versa) to a balanced ($|z^{+2}|\sim|z^{-2}|$) turbulent state could as well be associated with the steepening of the spectra from the low frequency $f^{-3/2}$ regime to the higher frequency $f^{-5/3}$ regime. A similar gradual change from an imbalanced towards a relatively balanced state is also evident with increasing heliospheric distance $R$. Even though $\sigma_{c}$ shows sufficiently higher values being associated with FSW, it declines slowly as understood from the straight line fit having a slope $\alpha=-0.05$ (see Fig. 6 bottom). This is again consistent with the absence of the $f^{-3/2}$ regime at $R=3.16$ au and recent observations of change in the inertial range spectral index from $-3/2$ to $-5/3$ with increasing $R$ (Chen et al., 2020; Shi, C. et al., 2021; Sioulas et al., 2023a).
We further determine the evolutionary nature of the break scale, $\tau_{K}$, with $R$ and have investigated its relationship with the typical ion and correlation scales. In Fig. 7 (top), we show the radial evolution of $\tau_{K}$, appearing in the scaling of $K$, converted from time scale to length scale ($l_{K}$) via Taylor’s hypothesis as mentioned in Section 3. Clearly, $l_{K}$ shift towards larger scales with $R$, as evident from Fig. 5 and Table 2 as well. We see that a strong power-law relation exists between $R$ and $l_{K}$, with $l_{K}$ evolving as $l_{K}\propto R^{\text{ }1.18}$ for $R<1$ au and $l_{K}\propto R^{\text{ }1.87}$ for $R>1$ au. The central panel in Fig. 7 shows how the break scale behaves with $R$ when normalized to the ion-inertial length scale, $d_{i}=c/\omega_{pi}$ (where $\omega_{pi}$$=\sqrt{ne^{2}/\epsilon_{0}m}$ is the plasma frequency). The ion-inertial scale has been found to vary between $\sim 45$ to $\sim 500$ km for $R$ ranging from $R\simeq$ 0.3—3.2 au. After normalization, we find that the evolutionary nature is nearly lost for FSW intervals in the inner heliosphere near the ecliptic plane, with a residual weak $R^{\text{ }0.13}$ dependence, and $l_{K}$ is $\sim 10^{3}$ times $d_{i}$. A similar pattern was observed (but not shown) after normalization with the ion gyro-radius $\rho_{i}=v_{th}^{\perp}/\Omega_{i}$, in the inner heliosphere (the $\rho_{i}$ in the outer heliosphere could not be computed again due to data limitations). Note that the typical ion scales have an approximately linear radial increase up to 5 au (e.g., see Bruno & Trenchi, 2014) which might explain the constant radial trend of the normalized break scale. However, beyond $1$ au, it is to be noted that even after normalization, the evolutionary nature of $l_{K}$ still persists, so that only the radial trend of the break decouples from that of the ion scales. The residual power law could be associated with the variation in heliospheric latitude (and to the associated variation of the angle between the large-scale magnetic field $B$ and the bulk speed $V_{sw}$) at which the FSW streams were sampled, indicated in the labels in Fig. 7. Understanding this variation of $l_{K}$ with latitude and $V_{sw}$-$B$ angle would be interesting for a future study but is currently beyond the scope of this paper. In order to compare the break scale $l_{K}$ and the correlation scale $L_{c}$, we have drawn $l_{K}$ normalized to $L_{c}$ as a function of $R$ (see Fig. 7 bottom). It is evident from the plot that, for $R<1$, a small power-law exponent is observed, $l_{K}/L_{c}\sim R^{\text{ }-0.14}$, so that the normalization to the correlation scale removes the radial dependence, similar to what we observe when normalized to the ion scale. Moreover, in this case, $l_{K}$ is $\sim 0.4$ times $L_{c}$ and certainly does not correspond to scales within the $f^{-1}$ power law regime in the spectrum, contrary to what has been suggested previously (Sorriso-Valvo et al., 2023). For $R>1$, $l_{K}$ approaches $L_{c}$, thereby explaining the absence of the $f^{-3/2}$ regime in the $R=3.16$ au interval and supporting recent observations of spectral steepening of the inertial range with increasing $R$ (Chen et al., 2020; Shi, C. et al., 2021; Sioulas et al., 2023a). Note that, in the inner heliosphere, break scales normalized to both the characteristic ion scale and the correlation scale follow weak radial dependence of $R^{\text{ }0.13}$ and $R^{\text{ }-0.14}$, respectively.
@@
Figure 8: Magnetic field power spectral trace plotted for nearly equidistant frequencies of the intervals S1 (blue), AS1 (green) and F6 (red) during solar maximum. The SSW and ASSW exhibits a broad inertial range while a $f^{-1}$ break is present in the case of FSW. In each range, a power law fit is shown along with the corresponding spectral exponent.
Figure 9: Kurtosis ($K$) of the magnetic field fluctuations as a function of time scale $\tau$ for several intervals of SSW (blue), ASSW (grren) and FSW (red), during solar maximum.

### v2

@@ 4.1 Observations during Solar Minimum
Figure 3: Magnetic power spectral trace of FSW intervals (smoothed using a running mean window) at varying heliospheric distances $R$ during solar minima. Top panels: Helios data (year 1976), bottom panels: Ulysses data (years 1995-1996). In all panels, the distance of the interval from the Sun is indicated. Vertical lines indicate the $f^{-1}$ break (dashed), the newly observed break $f_{b}$ (dot-dashed, separating the light and deep blue areas jointly forming the traditional inertial range), and the ion-scale break (Ulysses only). In each range, a power-law fit is shown (coloured lines) along with the corresponding spectral exponent.
Figure 4: Kurtosis $K(\tau)$ of magnetic field fluctuations for several intervals of FSW during periods of solar minima. Top panels: Helios data in the inner heliosphere from a sustained coronal hole. Bottom panels: Ulysses data in the outer heliosphere at varying distances and latitudes (the distance of each interval is indicated and associated with a given colour). Power-law fits and the corresponding scaling exponents are indicated. Vertical lines indicate the observed break, $\tau_{K}$ (solid lines), and the timescale corresponding to the spectral break, $1/f_{b}$ (dashed).
@@
During a period of solar minimum in 1976, using data from Helios spacecraft, we study FSW streams in the inner heliosphere (at $0.3,0.41,0.65,$ and $0.98$ au) from a sustained coronal hole near the ecliptic plane. Beyond $1$ au, FSW streams are studied using Ulysses data collected during the 1995-1996 solar minimum at varying heliospheric distances (at $1.44,2.1,2.75,$ and $3.16$ au), which were also measured at different latitudes. In Fig. 3, we have drawn the magnetic power spectral traces, smoothed using a running mean window. Top panels refer to Helios intervals, while bottom panels to Ulysses. As typically observed in the Alfvénic solar wind, at low frequencies we can identify a large-scale, energy-containing range (white background in the figure), where the power decays as $\sim f^{-1}$. Fitted power laws and the corresponding scaling exponents are shown as green lines. A break identifies a clear change in the power-law scaling exponent, as indicated by vertical dashed lines. Such break can be associated with the correlation scale of the turbulence. The low-frequency range is clearly visible in Helios data, while it is only indicatively present in the Ulysses intervals. This is consistent with the well-known shift of the correlation scale towards lower frequency with increasing $R$ in the solar wind (Davis et al., 2023). The $f^{-1}$ range is followed by the usual inertial range of turbulence, where the spectrum roughly follows an $f^{-5/3}$ power law dependence (Bruno & Carbone, 2005, 2013). However, a more accurate inspection shows that a further break emerges within such range, indicated by the vertical dot-dashed lines separating the light and deeper blue shaded areas in Fig. 3. Although the dynamical range of frequencies is relatively short, for intervals other than that at $3.16$ au, it is possible to identify two different sub-ranges with different power laws as demonstrated by the red and blue lines, with the associated scaling exponents indicated nearby. In the lower-frequency range (light blue background), the spectral index approaches ${-3/2}$, whereas at larger frequencies (deep blue background), the spectra show a transition to a ${-5/3}$ spectral index usually observed in non-Alfvénic solar wind (Bruno & Carbone, 2005, 2013; Alexandrova et al., 2009; D’Amicis et al., 2018). In isotropic turbulence, whereas an $f^{-5/3}$ scaling often represents an energy cascade by eddy fragmentation in strong turbulence, $f^{-3/2}$ scaling can possibly be explained by an energy cascade through the sporadic interaction of Alfvénic wave packets in MHD turbulence (Kolmogorov, 1941; Iroshnikov, 1963; Kraichnan, 1965). However, $-5/3$ and $-3/2$ power laws can also be obtained under various circumstances if anisotropy is taken into account (Goldreich & Sridhar, 1995, 1997; Boldyrev, 2006; Chandran et al., 2015). Irrespective of the true nature of energy cascade, a single power law is often assumed for the magnetic power spectra in the frequency range $10^{-4}-10^{-1}$ Hz (Bruno & Carbone, 2005, 2013), although a few studies have found variation in the power law exponents in the inertial range of magnetic power spectra (Wicks et al., 2011; Sioulas et al., 2023b; D’Amicis et al., 2025) as well as the scaling of higher order structure functions (Wu et al., 2022; Sorriso-Valvo et al., 2023; D’Amicis et al., 2025). In our study, the co-existence of the two sub-regimes (with $-3/2$ and $-5/3$ spectral indices) within the turbulence spectra of FSW has been consistently observed at various heliospheric distances both in the inner as well as the outer heliosphere. The break scale between those two sub-ranges, $f_{b}$, appears to shift towards lower frequencies (approaching the correlation scale) with increasing heliospheric distance. This is consistent with the fact that, when using single power law, a ${-3/2}$ scaling has been observed for solar wind close to the Sun, whereas a steeper ${-5/3}$ power law is obtained at and beyond $1$ au (Chen et al., 2020; Shi, C. et al., 2021; Sioulas et al., 2023a). Finally, in the Ulysses intervals, the ion-scale breaks are visible, separating the MHD range from the sub-ion range, where Hall effects and other kinetic effects start to affect the cascade (white background) (Banerjee & Galtier, 2016; Halder et al., 2023). Such break is usually observed at frequencies $\sim 10^{-1}$ Hz, which is the upper cut-off for the MHD range. However, similar breaks do not turn up in the Helios intervals due to the low cadence of the data used here.
@@
To further investigate on the sub-inertial range spectral break, $f_{b}$, we study the kurtosis $K(\tau)$ for all of the eight FSW intervals. The scaling of $K(\tau)$ defined in Section 3 are depicted for Helios and Ulysses data in Fig. 4 top and bottom panels, respectively, for each $R$. To inspect on the general radial trend of intermittency, we have drawn a consolidated plot for the Helios and Ulysses intervals (see Fig. 5). From this figure one can conclude that the value of $K$ at all scales increases with increasing $R$, thus implying higher intermittency with increasing heliospheric distance, in agreement with previous studies (Bruno et al., 2003; Sorriso-Valvo et al., 2023; Sioulas et al., 2023a). At each given distance $R$, $K$ is systematically found to decrease as one moves towards the larger scales. This is consistent with the notion that deviation from Gaussian statistics increases at smaller scales (Frisch, 1995; Sorriso-Valvo et al., 1999). Upon reaching the typical correlation scales of the flow ($\tau\simeq 10^{4}$ s), corresponding to the $f^{-1}$ power law in energy spectrum (see Fig. 3), the kurtosis saturates to a constant value $K\simeq 1.67$, representing a quasi-Gaussian distribution (with a non-zero skewness) of the fluctuations of the magnetic field components (see Appendix A). Within the inertial range, from the nature of $K(\tau)$ in Fig. 4, a clear signature of broken power law is observed. While two breaks are visible for Ulysses data (with $1$ s resolution), the small-scale break at around $\tau\sim 10$ s is missing for the intervals using Helios magnetic field data with $6$ s resolution. The other break which occurs at a larger $\tau$ (solid vertical lines) is clearly visible for both Helios and Ulysses data. As observed for the spectra, this break scale, $\tau_{K}$, shifts towards larger scales as $R$ increases. Within the distance range of 0.30.30.30.32.752.752.752.75 au, $\tau_{K}$ is found to increase from $\sim 100$ s to $\sim 1500$ s. It is to be emphasized here that except for certain cases, the appearance of the break $\tau_{K}$ is persistent in the component-wise $K$ scaling as well (see Figs. 10 and 11 in Appendix B). A detailed list of the break scale $\tau_{K}$ as a function of $R$ is given in Table 2. As it is evident from Fig.s 4 and 5, $\tau_{K}$ separates the steeper power law ($K\sim\tau^{-\kappa}$ with $\kappa\simeq 0.37$ averaged over the eight intervals) at smaller scales (dashed lines) from the less steeper one ($\kappa\simeq 0.11$ on average) at large scales (dotted lines), but with an exception. Note that for the Ulysses interval at $R=3.16$ au, $K(\tau)$ reaches the Gaussian regime without going through the large-scale break, suggesting that the turbulence has fully developed, which transforms the shallower scaling range at large scale into the steeper power law at smaller scales.
@@
As mentioned in the introduction, similar broken power-law behaviour for $K(\tau)$ in FSW has already been observed by Sorriso-Valvo et al. (2023). However, those authors suggested that $\tau_{K}$ might correspond to the break between low-frequency $f^{-1}$ regime to Kolmogorov $f^{-5/3}$ regime in the magnetic power spectra. This was inspired by the fact that $f^{-1}$ regime is exclusively found in FSW intervals, and the $f^{-1}$ break also shows nearly similar behaviour to $1/\tau_{K}$ as $R$ changes (Davis et al., 2023). Instead, for all the intervals where the break is observed, it is systematically found in our study that $1/\tau_{K}$ occurs at a higher frequency (roughly by a factor $10$) than the $f^{-1}$ break scale (see Fig. 3). The inverse of $\tau_{K}$ is typically corresponding to $f_{b}$, although with some consistent small discrepancy that could be due to the different frequency response of Fourier transform and scale-dependent increments (see Fig. 4 where both $\tau_{K}$, solid lines, and $1/f_{b}$, dashed lines, are drawn). The two scaling ranges in the kurtosis therefore approximately correspond to the two inertial sub-ranges observed in the spectrum. Since $PSD$ and kurtosis are related quantities, the observation of a double power law in both supports the robustness of the break and, therefore indicates the emergence of a new characteristic scale in the inertial range that marks the transition from $f^{-3/2}$ to $f^{-5/3}$ regime.
@@
Figure 6: Top: Normalized cross-helicity spectrum $\hat{\sigma_{c}}$ (smoothed using a running mean window) of the four Helios FSW intervals in the inner heliosphere. The light and deep blue shaded regions depict the $-3/2$ and $-5/3$ regimes respectively. The spectral break frequency $f_{b}$ and the frequency associated with the kurtosis break, $\tau_{K}^{-1}$, are indicated by dot-dashed and dashed lines respectively. Bottom: Cross-helicity $\sigma_{c}$ of all the FSW intervals as a function of the heliospheric distance $R$.
@@
Figure 7: FSW during solar minima. Top: Kurtosis scaling break $l_{K}(=V_{sw}\tau_{K})$ versus the heliospheric distance $R$. The same break scale $l_{K}$ normalized by the ion-inertial length scale $d_{i}$, $l_{K}/d_{i}$ (middle) and $l_{K}$ normalized by the correlation length $L_{c}$ (bottom) as a function of $R$. The different colours reproduce the colours in Fig. 3, and in the Ulysses intervals, the latitude is indicated. Two power laws were identified in the inner and outer heliosphere, respectively. The fitted power laws and the corresponding parameters are indicated.
@@
Figure 8: Magnetic field power spectral trace (smoothed using a running mean window) of the FSW, ASSW & SSW intervals tabulated in Table-1 during solar maximum (the PSD’s are artificially shifted for representation). The SSW exhibits a broad inertial range while different power law breaks are present in the case of ASSW & FSW. In each range, the compensated spectra is shown along with the corresponding spectral exponent and the normalized cross-helicity.
[…]
We further determine the evolutionary nature of the break scale with radial distance, and investigate its relationship with the typical ion and correlation scales. In Fig. 7 (top), we show the radial evolution of $\tau_{K}$, appearing in the scaling of $K$, converted from time scale to length scale, $l_{K}$, via Taylor’s hypothesis as mentioned in Section 3. Clearly, $l_{K}$ shift towards larger scales with $R$, as evident from Fig. 5 and Table 2 as well. We see that a robust power-law relation exists between $R$ and $l_{K}$, with $l_{K}$ evolving as $l_{K}\propto R^{\text{ }1.18}$ for $R<1$ au and $l_{K}\propto R^{\text{ }1.87}$ for $R>1$ au. The central panel in Fig. 7 shows how the break scale behaves with $R$ when normalized to the ion-inertial length scale, $d_{i}=c/\omega_{pi}$ (where $\omega_{pi}$$=\sqrt{ne^{2}/\epsilon_{0}m}$ is the plasma frequency). The ion-inertial scale has been found to vary between $\sim 45$ to $\sim 500$ km for $R$ ranging from $R\simeq$ 0.3—3.2 au. After normalization, we find that the evolutionary nature is nearly lost for FSW intervals in the inner heliosphere near the ecliptic plane, with a residual weak $R^{\text{ }0.13}$ dependence, and $l_{K}$ is $\sim 10^{3}$ times $d_{i}$. A similar pattern was observed (but not shown) after normalization with the ion gyro-radius $\rho_{i}=v_{th}^{\perp}/\Omega_{i}$, in the inner heliosphere (the $\rho_{i}$ in the outer heliosphere could not be computed again due to data limitations). Note that the typical ion scales have an approximately linear radial increase up to 5 au (e.g., see Bruno & Trenchi, 2014), which might explain the constant radial trend of the normalized break scale. However, beyond $1$ au, it is to be noted that even after normalization, the evolutionary nature of $l_{K}$ still persists so that only the radial trend of the break decouples from that of the ion scales. The residual power law could be associated with the variation in heliospheric latitude (and to the associated variation of the angle between the large-scale magnetic field $B$ and the bulk speed $V_{sw}$) at which the FSW streams were sampled, indicated in the labels in Fig. 7. Understanding this variation of $l_{K}$ with latitude and the angle between the bulk wind speed and the large-scale magnetic field would be interesting for a future study but is currently beyond the scope of this paper. In order to compare the break scale $l_{K}$ and the correlation scale $L_{c}$, we have drawn $l_{K}$ normalized to $L_{c}$ as a function of $R$ (see Fig. 7 bottom). Here $L_{c}$ is the Taylor-shifted $\tau_{c}$, which is the time lag at which the trace of the correlation matrix of $\mathbf{B}$ decreases to $1/e$ of its initial value. It is evident from the plot that, for $R<1$, a small power-law exponent is observed, $l_{K}/L_{c}\sim R^{\text{ }-0.14}$, so that the normalization to the correlation scale removes the radial dependence, similar to what we observe when normalized to the ion scale. Moreover, in this case, $l_{K}$ is $\sim 0.4$ times $L_{c}$ and certainly does not correspond to scales within the $f^{-1}$ power law regime in the spectrum, contrary to what has been suggested previously (Sorriso-Valvo et al., 2023). For $R>1$, $l_{K}$ approaches $L_{c}$, thereby explaining the absence of the $f^{-3/2}$ regime in the $R=3.16$ au interval and supporting recent observations of spectral steepening of the inertial range with increasing $R$ (Chen et al., 2020; Shi, C. et al., 2021; Sioulas et al., 2023a). Note that, in the inner heliosphere, break scales normalized to both the characteristic ion scale and the correlation scale follow a comparably weak radial dependence of $R^{\text{ }0.13}$ and $R^{\text{ }-0.14}$, respectively.
Figure 9: Kurtosis ($K$) of the magnetic field fluctuations as a function of time scale ($\tau$) for several intervals of (a) FSW (orange), (b) ASSW (green) and (c) SSW (deep blue), during solar maximum. The normalized cross-helicity is also indicated in the legend. In panels (a), (b) and (c), the kurtosis for each interval have been artificially shifted for better representation while panel (d) shows a consolidated plot of all the intervals (with no artificial shifting) of the three types of solar wind.